OpenStudy (anonymous):

I NEED HELP. use the fact that the trigonometric functions are periodic to find the exact value of each expression. cos 420°

OpenStudy (anonymous):

periodic means that the values actually repeat themselves every period. (bad definition on my part lol) but, cos(420) = cos(420-360) = cos(60)

OpenStudy (anonymous):

how did you got 360

OpenStudy (anonymous):

@alexandercpark how did you got 360